OpenStudy (anonymous):

f(x) = sqrt x-5 find f-1

OpenStudy (anonymous):

@Marv93w

OpenStudy (anonymous):

x^2+5

OpenStudy (anonymous):

A. f -1 (x) = x2 + 5; y greater than or equal to 5 B. f-1 (x) = x2 +25; y greater than or equal to 5 C. f -1 (x) = x2 + 5; y greater than or equal to 5 D. f-1 (x) = x2 +25; y greater than or equal to 25

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

A and C look the same

OpenStudy (anonymous):

Yea ik i guess c is supposed to be greater than or equal to 25

OpenStudy (anonymous):

it y greater than or equal to 5

OpenStudy (anonymous):

ok